OpenStudy (anonymous):

Corner is the treasurer oh his school. He has to decide whether or not all of the activities fit into his budget. His budget is $200 without any fundraisers. He decided to estimate the cost of the activities. He decided to budget $14 for canned goods and $30 for school clean up day. Then he decided to split the remaining money evenly between the pep rally and the float. Write a two step equation to find how much money is budget for the pep rally. How much money is budgeted for the float?

OpenStudy (anonymous):

I'm not sure what a 'two step equation' is but this should give you all the info you need I think: x:= money after canned goods and cleanup y:= money for pep rally x=200-14-30=156 y= x/2=156/2=78 so there's $78 for both the pep rally and the float

OpenStudy (anonymous):

thank you